Electric Current

2 long, straight, parallel conductors, separated by 10cm, carry a current of 5A each. Calculate the magnetic induction at a point midway between the conductors when the currents are (i) in the same direction and (ii) in opposite directions.

Please Reply Soon.

Thanks.

Arvind Petkar
15 Years agoGrade

2 Answers

T.MANI TEJA
15 Years ago

1.whn the point lies between both the conductors carrying current in the same direction then the magnetic field due to left wire is in inward direction...while the magnetic field due to other wire is in upward direction....there fore total magnetic field is zero....

2.whn the currents in both the wires are in opposite direction then magnetic field due to both the condutors is either in upward direction or downward direction...there fore calculate it using formula....

vikas askiitian expert
ApprovedApproved Tutor Answer15 Years ago

i) current in same direction

magnetic field due to one wire is cancelled by other wire so net magnetic field becomes 0 ..

 

ii)current in opposite direction

magnetic field due to both wires at center is added , let B is magnetic field due to one wire

at center then total magnetic field is 2B ...

B = uoI/2pid

total magnetic field = 2B = uoI/pid

d = 10/2 = 5cm , I = 5A ,so

total magnetic field = 4x10-5 tesla

 

approve if u like my ans